Calculate the volume in liters of a 0.0026M calcium sulfate solution that contains 175.g of calcium sulfate. Round your answer to significant digits.
Answer #1

Answer

moles of CaSO4 = mass/molar mass= 175 g/136.14= 1.285 moles

Molarity= moles of CaSO4/volume of solution in L

volume of solution in L = moles of CaSO4/molarity

volume of solution in L = 1.285/0.0026 = 494 L